OpenStudy (anonymous):

7 How fast must a ball be rolled along the surface of a 70-cm high table so that when it rolls off the edge it will strike the floor at the same distance (70cm) from the point directly below the edge of the table?

OpenStudy (anonymous):

horizontal range = initial velocity multiplied by the time of descent.

OpenStudy (anonymous):

@sambex do you get it?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

v_y=g*t v_x=v_0 y_0=.7 x_0=0 x=v_0t y=y_0-(g*t^2)/2=0 t^2=2*y_0/g t=sqrt(2*y_0/g)=sqrt(1.4/9.81) x=v_0t=.7 v_0=.7/t=.7*sqrt(9.81/1.4)=1.853 m/s

OpenStudy (anonymous):

Do you understand each of the equations used to get to the answer and how/why I used each of them?

OpenStudy (anonymous):

@sambex